# EntroLLM: Entropy Encoded Weight Compression for Efficient Large Language Model Inference on Edge Devices

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate exceptional performance across various tasks, but their large storage and computational requirements constrain their deployment on edge devices. To address this, we propose EntroLLM, a novel compression framework that integrates mixed quantization with entropy coding to reduce storage overhead while maintaining model accuracy. Our method applies a layer-wise mixed quantization scheme – choosing between symmetric and asymmetric quantization based on individual layer weight distributions – to optimize compressibility. We then employ Huffman encoding for lossless compression of the quantized weights, significantly reducing memory bandwidth requirements. Furthermore, we introduce parallel Huffman decoding, which enables efficient retrieval of encoded weights during inference, ensuring minimal latency impact. Our experiments on edge-compatible LLMs, including smolLM-1.7B-Instruct, phi3-mini-4k-Instruct, and mistral-7B-Instruct, demonstrate that EntroLLM achieves up to $30\%$ storage reduction compared to uint8 models and up to $65\%$ storage reduction compared to uint4 models, while preserving perplexity and accuracy, on language benchmark tasks. We further show that our method enables $31.9\%$$146.6\%$ faster inference throughput on memory-bandwidth-limited edge devices, such as NVIDIA Jetson P3450, by reducing the required data movement. The proposed approach requires no additional re-training and is fully compatible with existing post-training quantization methods, making it a practical solution for edge LLMs.
Index Terms:
@@
Large Language Models (LLMs), quantization, entropy coding, Huffman coding, parallel decoding
I Introduction
@@
Large language models (LLMs) have demonstrated remarkable performance in various domains [1, 2], but their substantial size poses challenges for deployment, especially on resource-constrained edge devices [3]. For example, even a smaller LLM such as mistral-7B-Instruct [4] requires over $14$ GB of memory with weights encoded in $16$-bit floats (fp16), exceeding the capacity of most edge devices. As a direct result of being massively parameter-heavy, the primary bottleneck in LLM inference, particularly for generative tasks, is memory bandwidth rather than computation [5]. This means that the speed of loading and storing parameters becomes the most critical constraint [6].
@@
Figure 1: Overall Schematic that demonstrates our edge-device inference scheme. A floating point model is first quantized using a mixed quantization scheme to maximize storage compressibility. The model after quantization (blue) is the model that we will finally use during inference. But when storing this model on edge devices (green), we compress the integer weights using entropy coding. To ensure fast inference, we come up with a parallel decoding scheme to revert to integer weights from variable length entropy-coded weights. Section III has a detailed description of the various components described here.
@@
Figure 2: A visual explanation of the different uniform quantization grids [7] for a bit-width of $8$. $s$ is the scaling factor, $z$ the zero-point. The floating-point grid is in black, and the integer quantized grid is in blue. In our work, we use either a symmetric unsigned or an asymmetric quantization scheme on each layer based on the individual layer’s weight distribution.
@@
Quantization, a technique that encodes the model parameters in lower precision formats, has effectively reduced the size of LLMs [8]. However, further compression without compromising performance remains challenging, especially for smaller LLMs ($\leq$ 10B parameters) designed for edge applications. To address this, we focus on reducing the storage overhead for deploying these small LLMs on low-power edge devices. We introduce an entropy-based encoding scheme to further compress these models without loss of information. Specifically, we leverage variable-length bit widths to encode the quantized weights, using techniques rooted in entropy coding to achieve lossless compression. During inference, the original weights are recovered with precision without additional storage overhead, ensuring that performance degradation is minimized to minor quantization effects. This approach, illustrated in Figure 1, employs Huffman coding due to its alignment with the Shannon entropy bounds, offering optimal compression for the quantized model parameters while maintaining the computational efficiency required for edge inference. We also use mixed asymmetric and symmetric unsigned quantization to induce better weight compression and implement a partitioning scheme for model weights so that parallel decoding is made possible during inference.
Our main contributions to this work are as follows –
  1. 1.
@@
    Mixed Quantization Scheme: We introduce a layer-specific quantization strategy that balances storage efficiency with model performance, ensuring optimal weight distribution across layers.
  2. 2.
@@
    Huffman Weight Encoding: We leverage entropy-based Huffman coding to further compress quantized weights, reducing storage overhead while preserving inference accuracy.
  3. 3.
@@
II Related Works
Before the era of LLMs, various compression techniques were developed to enhance the efficiency of deep learning models. These techniques include pruning [11], quantization [12], knowledge distillation [13], and alternative numerical representations such as logarithmic numbers [14] and posit [15]. Although effective, most of these methods rely on full or partial retraining of the models to achieve optimal results. However, for LLMs, retraining is highly impractical, as their memory requirements during training exceed the capacity of most consumer and commercial GPUs available today. To bypass this challenge, recent research has proposed compression methods for LLMs that require minimal to no retraining.
II-A Pruning
@@
Sun et al. [16] proposed an approach that prunes weights with the smallest magnitudes multiplied by the corresponding input activations per output. LLM-pruner [17] adopted structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the major functionality of the LLM.
II-B Quantization
Numerous weight-only post-training quantization methods have been introduced, such as GPTQ [18], AWQ [19], and SpQR [20], to enhance the memory bandwidth of LLMs. To minimize accuracy loss, these techniques typically leverage a small calibration dataset to fine-tune the quantized weights. The primary goal of these approaches is to maintain the layer-wise output activations consistent with their full-precision counterparts, allowing weight precision to be reduced to $4$ bits with minimal performance impact. Although these works focus on uniform quantization and are commonly used, others focus on non-uniform quantization based on sensitivity [21, 5]. These works focus on mitigating the effects of parsimony on more sensitive weight values, allowing more aggressive quantization beyond $8$ bits without compromising performance. Recent studies have also explored ultra-low-precision models [22, 23], such as ternary LLMs [24], which address outliers in both activations and weights. However, the applicability of these advanced quantization techniques to smaller LLMs designed for edge devices remains largely unexplored.
II-C Codebook-based Entropy Coding
Codebook-based entropy coding [25] compresses model parameters by mapping frequent symbols to shorter codes using predefined dictionaries. Although this method is effective in reducing encoding length and offers sub-1-bit compression, it is primarily suited for mixture-of-experts models and lacks a practical solution for resource-constrained edge devices. Moreover, codebook-based compression is not Shannon-rate optimal, which limits its efficiency. In contrast, our approach employs Huffman encoding, which achieves optimal compression rates, and we further enhance inference efficiency by implementing parallel Huffman decoding to meet the stringent performance demands of edge hardware inference.
III Proposed Method
Entropy coding aims to represent the data in the most compact way possible by leveraging the frequency of symbols in the source material. This approach is grounded in Shannon’s entropy, which measures the average amount of information produced by a stochastic data source. Huffman encoding [26], developed by David A. Huffman in 1952, optimally achieves this by assigning shorter codes to more frequent symbols and longer codes to less frequent ones, ensuring that the total size of the encoded data is minimized. In our approach, Huffman encoding is our main workhorse for achieving beyond-quantization on-device compression.
III-A Mixed Quantization Scheme
In this work, we aim to compress the weights further beyond post-training quantization. The weights of a neural network are normally distributed [27] after training, and after quantization, their distribution should remain unchanged. To increase compressibility from entropy coding, we need to ensure that the quantized weights are distributed in a way that minimizes entropy. To achieve this, we employ a mixed quantization scheme that quantizes the parameters of each layer with symmetric unsigned quantization (equation 1) or asymmetric quantization (equation 2), depending on the weight distribution of that layer. The different quantization schemes and how they map from floating to integer grids are explained in figure 2.
$\mathtt{W_{int}}$ $\leftarrow\left\lfloor\frac{\mathtt{W_{fp}}}{s}\right\rceil$ (1)
$\mathtt{W_{int}}$ $\leftarrow\left\lfloor\frac{(\mathtt{W_{fp}}-z)}{s}\right\rceil$ (2)
@@ III-A Mixed Quantization Scheme
Figure 3: In the serial decoding setting, parallelization is not possible as the variable code lengths make it difficult for us to predict the start token of a symbol in our remaining encoded parameters. By keeping the parameter space’s original weight tensor packing structure intact, we can parallelize decoding as now we can assign different encoded tensors (chunks), to individual threads assigned to cores. Modern LLMs have hundreds to thousands of such weight tensors, and hence, coarse-grained parallelism is achievable. In the figure <p> represents current bits being decoded, while <unk> represents unknown encoded parameter bits.
Algorithm 1 Entropy-based LLM Compression and Inference
1:  Input: Pre-trained LLM with floating-point weights $\mathtt{W_{fp}}$
2:  Output: Compressed model for efficient edge deployment
3:  procedure Cloud Processing $(\mathcal{C})$
4:      for each layer $\mathtt{k}$ in the model $\triangleright$ Quantization
5:          if $\mathtt{W_{fp}^{k}|_{max}\times W_{fp}^{k}|_{min}\geq 0}$ then
6:              $\mathtt{W_{int}^{k}}\leftarrow\left\lfloor\frac{\mathtt{W_{fp}^{k}}}{s}\right\rceil$
7:          else
8:              $\mathtt{W_{int}^{k}}\leftarrow\left\lfloor\frac{(\mathtt{W_{fp}^{k}}-z)}{s}\right\rceil$    $\triangleright$ $z$ is zero-point, $s$ is scaling factor
9:          end if
10:      end for
11:      $F\leftarrow\mathcal{F}\{\mathtt{W_{int}^{0}},\mathtt{W_{int}^{1}},\dots,\mathtt{W_{int}^{k}},\dots\}$$\triangleright$ Compute    frequency of each quantized weight value across model
12:      $H,P\leftarrow\mathcal{H}\{F\}$$\triangleright$ Construct Huffman tree to generate    symbol lookup table and probability statistics
13:      for each layer $\mathtt{k}$ in the model do
14:          $\mathtt{W_{c}^{k}}\leftarrow H\{\mathtt{W_{int}^{k}}\}$
15:      end for
16:      Store model metadata: $H,P,\{\mathtt{W_{c}}\}^{\mathtt{k}}$     $\triangleright$ Preserving    the weight tensor packing structure $\{\mathtt{W_{c}}\}^{\mathtt{k}}$    $\triangleright$ Enables parallel decoding during inference
17:  end procedure
18:  procedure Edge Device Operations $(\mathcal{D})$
19:      Load: $H,P,\{\mathtt{W_{c}}\}^{\mathtt{k}}$
20:      Initialize: parallel decoding with $T$ threads
21:      for each thread $t\in T$ in parallel do
22:      Assign: $\{\mathtt{W_{c}}\}^{\mathtt{m}}$ to thread $t\in T$ to balance workload
23:          for each assigned layer $\mathtt{m}$ do
24:              $\mathtt{W_{int}^{m}}\leftarrow H^{-1}\{\mathtt{W_{c}^{m}}\}$
25:              Perform layer computation using $\mathtt{W_{int}^{m}}$
26:          end for
27:      end for
28:  end procedure
This dual quantization scheme ensures that each layer retains its inherent distribution, enabling the network to operate within a consistent and normalized range across layers. By carefully choosing the quantization method based on the sign characteristics of each layer, we normalize the parameter distributions, resulting in a system in which all layers, regardless of their original distribution, contribute to a final, highly Gaussian distribution for all parameters in the model. This convergence of distributions is the key to our subsequent Huffman encoding step. Using the low entropy produced by combining these balanced weight distributions, we generate an efficient Huffman tree, allowing us to encode the model weights with minimal redundancy while maintaining precision during inference. This synergy between the quantization strategy and entropy coding not only compresses the model effectively but also ensures that the performance remains unaffected by the transformation. Lines 4 to 10 in algorithm 1 describe the quantization scheme used.
@@
Figure 4: Top: Parameter distribution for 8-bit quantized models. Bottom: Parameter distribution for 4-bit quantized models. Model size increases from left to right.
@@
Effective Bits $16$ $\mathbf{5.92}$ $\mathbf{1.57}$ $16$ $\mathbf{5.58}$ $\mathbf{1.39}$ $16$ $\mathbf{5.84}$ $\mathbf{1.62}$
Wikitext2 (ppl.) $\downarrow$ $\mathbf{23.81}$ $23.93$ $24.14$ $\mathbf{9.03}$ $9.44$ $10.10$ $\mathbf{8.17}$ $8.24$ $8.29$
HellaSwag (acc.) $\uparrow$ 25.85% $25.55\%$ $25.30\%$ 82.2% $82.10\%$ $81.01\%$ 58.37% $58.33\%$ 58.21%
GSM8k CoT (acc.) $\uparrow$ - - - 77.37% $72.84\%$ $70.58\%$ 52.2% $48.62\%$ $45.36\%$
@@
TABLE I: Benchmarks: Perplexity & Accuracy benchmarks for smolLM-1.7B-Instruct, phi3-mini-4k-Instruct and mistral-7B-Instruct on various language tasks
III-B Huffman weight encoding
@@
Huffman encoding – which originates from entropy coding – is a popular data compression technique. The technique is based on the construction of a binary tree, where each symbol is assigned a score corresponding to its frequency of occurrence. This process involves recursively combining pairs of symbols into a tree structure. More frequent symbols are positioned closer to the root of the tree, which results in shorter paths and, consequently, shorter binary codes for these symbols. The length of each symbol’s binary code is proportional to its distance from the root, thereby minimizing the total length of the encoded data. This ensures that no other prefix code can offer a more efficient representation of the data, making the Huffman encoding optimal for lossless compression.
In this work, we employ Huffman encoding to compress the weights of quantized LLMs, enabling their deployment on edge devices. By applying Huffman encoding to the quantized weights of an LLM, particularly those with skewed value distributions, we achieve significant compression without sacrificing accuracy. As discussed in Section II-B, a post-training quantized LLM can match the performance of a full-precision LLM for most language generation benchmarks. Huffman coding converts the model’s weight parameters into variable-length codes, assigning shorter codes to frequent weight values, which reduces both storage and bandwidth costs. Lines 11 to 16 in algorithm 1 describe the Huffman encoding of weights post-quantization and how the encoded weights are stored to ensure on-the-fly parallel decoding during inference.
III-C Parameter space segmentation for parallel Huffman decoding
To achieve efficient real-time inference with Huffman-encoded model weights, we introduce a method for parallelizing the decoding process. Huffman decoding is inherently sequential, as it involves interpreting variable-length encoded symbols one at a time. This sequential nature increases the latency of the LLMs when deployed on edge devices, limiting their potential for real-time execution. To overcome this limitation, we propose a segmentation strategy that preserves the original structure of the encoded weights’ parameter space during Huffman encoding.
The encoded parameter space is divided into discrete segments, maintaining the inherent organization of the weight tensors. This preservation of the structure allows for each segment to be decoded independently. Each of these segments can contain an average number of symbols and can be decoded in parallel, allowing concurrent decoding across multiple threads. By preserving the weight-tensor packing structure, we ensure that the beginning and end of each encoded segment are known in advance. This enables independent decoding without the need for synchronization between threads. By doing this, we can distribute the segments across multiple threads, allowing them to be decoded in parallel. The design ensures that each thread operates on its portion of the encoded data without having to wait for others to complete, thus bypassing the serial bottleneck that typically hinders Huffman decoding.
An additional challenge arises from the fact that individual segments can exhibit skewed distributions of Huffman-encoded symbols, leading to potential imbalances in decoding times. To address this, we employ a shuffling mechanism in which multiple segments are assigned to each thread. This ensures that, on average, each thread processes a balanced workload, as the mixture of segments helps to even out any variance in decoding complexity caused by the skewed distributions of longer or shorter symbols. This approach not only accelerates the overall decoding process but also guarantees that the system is well-suited for real-time inference scenarios, where maintaining low latency is paramount. By parallelizing the decoding of Huffman-encoded weights while preserving the original parameter space structure, we achieve both efficient model restoration and the rapid inference required in time-critical applications. The schematic for parallel decoding is represented diagrammatically in figure 3. The algorithm 1 demonstrates the proposed methodology end-to-end as described in sections III-A, III-B, and III-C.
IV Experiments

### v4

@@ Abstract
Large Language Models (LLMs) achieve strong performance across tasks, but face storage and compute challenges on edge devices. We propose EntroLLM, a compression framework combining mixed quantization and entropy coding to reduce storage while preserving accuracy. We use a combination of unsigned and asymmetric quantization. Tensor-level quantization produces an entropy-reducing effect, increasing weight compressibility, and improving downstream Huffman encoding by $7\times$ (8-bit) and $11.3\times$ (4-bit) over state-of-the-art methods. Huffman coding further reduces memory bandwidth demands, while a parallel decoding strategy enables efficient weight retrieval with minimal latency. Experiments on edge-scale LLMs (smolLM-1.7B, phi3-mini-4k, mistral-7B) show up to 30% storage savings over uint8 and 65% over uint4 models, with 31.9–146.6% faster inference on memory-limited devices like the NVIDIA Jetson P3450. EntroLLM requires no retraining and is compatible with existing post-training quantization pipelines, making it practical for edge LLM deployment.
@@
Index Terms—  Large Language Models (LLMs), quantization, entropy coding, Huffman coding, parallel decoding
1 Introduction
@@
Large language models (LLMs) have demonstrated remarkable performance in various domains [5, 29], but their substantial size poses challenges for deployment, especially on resource-constrained edge devices [32]. For example, even a smaller LLM such as mistral-7B-Instruct [18] requires more than $14$ GB of memory with weights encoded in $16$-bit floats (fp16), exceeding the capacity of most edge devices. As a direct result of being massively parameter-heavy, the primary bottleneck in LLM inference, particularly for generative tasks, is memory bandwidth rather than computation [19]. This means that the speed of loading and storing parameters becomes the most critical constraint [12].
@@
Fig. 1: Overall schematic of our edge-device inference scheme. A floating-point model is first quantized using a mixed scheme to maximize compressibility. For storage on edge devices, integer weights are entropy-coded, then decoded in parallel during inference to minimize latency. Section 3 details these components.
(a) Symmetric Signed
(b) Unsigned
(c) Asymmetric
@@
Fig. 2: A visual explanation of the different uniform quantization grids [23] for a bit-width of 88. $s$ is the scaling factor, $z$ the zero-point. The floating-point grid is in black, and the integer quantized grid is in blue. In our work, we use either an unsigned or an asymmetric quantization scheme on each layer based on the individual layer’s weight distribution.
@@
Quantization, a technique that encodes the model parameters in lower precision formats, has effectively reduced the size of LLMs [31]. However, further compression without compromising performance remains challenging, especially for smaller LLMs ($\leq$ 10B parameters) designed for edge applications. To address this, we focus on reducing the storage overhead for deploying these small LLMs on low-power edge devices. We introduce an entropy-based encoding scheme to further compress these models without loss of information. Specifically, we leverage variable-length bit widths to encode quantized weights, using techniques rooted in entropy coding to achieve lossless compression. During inference, the original weights are recovered with precision without additional storage overhead, ensuring that performance degradation is minimized to minor quantization effects. This approach, illustrated in Figure 1, employs Huffman coding due to its alignment with the Shannon entropy bounds, offering optimal compression for the quantized model parameters while maintaining the computational efficiency required for edge inference. We also use a tensor-level mixed quantization scheme, comprising asymmetric quantization and unsigned quantization to induce better weight compression, and implement a partitioning scheme for model weights so that parallel decoding is made possible during inference. Our main contributions to this work are as follows –
@@
  • Mixed Quantization Scheme: We introduce a layer-specific quantization strategy that balances storage efficiency with model performance. This approach significantly improves the downstream compressibility of entropy compared to state-of-the-art quantization methods (as shown in table 2).
@@
  • Huffman Weight Encoding: We leverage entropy-based Huffman coding to compress quantized weights further, reducing storage overhead while preserving accuracy.
@@
2 Related Works
Before LLMs, compression techniques such as pruning [13], quantization [16, 3], knowledge distillation [14], and alternative number formats such as logarithmic [25] and posit [6] were developed to improve deep learning efficiency. These typically require retraining, which is impractical for LLMs due to extreme memory demands. Recent work focuses instead on LLM compression with minimal or no retraining.
@@ 2 Related Works
Pruning: Sun et al. [28] proposed an approach that prunes weights with the smallest magnitudes multiplied by the corresponding input activations per output. LLM-pruner [22] adopted structural pruning that selectively removes non-critical coupled structures based on gradient information, maximally preserving the main functionality of the LLM.
Quantization: Post-training quantization methods such as GPTQ [11], AWQ [21], and SpQR [9] reduce LLM memory by fine-tuning weights with small calibration datasets, allowing $4$-bit precision with minimal accuracy loss. Although most focus on uniform quantization, sensitivity-based non-uniform approaches [8, 19] allow for more aggressive quantization. Ultra-low-precision models, including ternary LLMs [4, 30, 7], address activation and weight outliers. However, their use for small edge-focused LLMs remains largely unexplored.
3 Proposed Method
Entropy coding compresses the data by exploiting symbol frequencies. Huffman encoding [17] optimally assigns shorter codes to frequent symbols and is the core of the beyond-quantization compression framework in our work.
Mixed Quantization Scheme: We aim to compress weights beyond post-training quantization. State-of-the-art models often utilize block-based quantization (e.g., blocks of 32 weights), which flattens the integer set and increases entropy. In contrast, obtaining a set of integer codes for an entire tensor natively results in a much “spikier” distribution with lower entropy. To preserve accuracy while leveraging this tensor-level compressibility, we employ a mixed quantization scheme, applying either unsigned (left of Eq. 1) or asymmetric (right of Eq. 1) quantization per layer, based on its weight distribution. Fig. 2 illustrates these schemes and their mapping from floating-point to integer grids.
$\mathtt{W_{int}}\leftarrow\left\lfloor\frac{\mathtt{W_{fp}}}{s}\right\rceil,\ \ \ \mathtt{W_{int}}\leftarrow\left\lfloor\frac{(\mathtt{W_{fp}}-z)}{s}\right\rceil$ (1)
@@ 3 Proposed Method
Fig. 3: In the serial decoding setting, parallelization is not possible as the variable code lengths make it difficult for us to predict the start token of a symbol in our remaining encoded parameters. By keeping the parameter space’s original weight tensor packing structure intact, we can parallelize decoding, as now we can assign different encoded tensors (chunks) to individual threads assigned to cores. Modern LLMs have hundreds to thousands of such weight tensors, and hence, coarse-grained parallelism is achievable. In the figure <p> represents current bits being decoded, while <unk> represents unknown encoded parameter bits.
By quantizing at the tensor level, we maintain a “spikier” distribution of integer codes compared to block-based approaches. This spikiness directly translates into lower entropy and higher compressibility. Using the low entropy produced by this tensor-level strategy, we generate an efficient Huffman tree, allowing us to encode the model weights with minimal redundancy while maintaining precision during inference. This synergy between the quantization strategy and entropy coding not only compresses the model effectively but also ensures that the performance remains unaffected by the transformation. As shown in Table 2, our quantization scheme improves downstream entropy compressibility by a factor of up to $\mathbf{8.1\times}$ for $8$-bit models and $\mathbf{13.1\times}$ for $4$-bit models compared to SOTA techniques.
@@
Fig. 4: Top: Parameter distribution for 8-bit quantized models. Bottom: Parameter distribution for 4-bit counterparts. Model size increases from left to right.
@@
(SOTA) (SOTA) (SOTA) (SOTA) (SOTA) (SOTA)
Wikitext2 (ppl.) $\downarrow$ $\mathbf{23.81}$ $23.92$ $24.15$ $\mathbf{9.03}$ $9.45$ $10.11$ $\mathbf{8.17}$ $8.22$ $8.28$
HellaSwag (acc.) $\uparrow$ 25.87% $25.55\%$ $25.28\%$ 82.2% $82.11\%$ $81.00\%$ 58.37% $58.34\%$ 58.23%
GSM8k CoT (acc.) $\uparrow$ - - - 77.37% $72.85\%$ $70.57\%$ 52.2% $48.63\%$ $45.38\%$
Weight Encoding - uint8 uint4 - uint8 uint4 - uint8 uint4
(Ours) (Ours) (Ours) (Ours) (Ours) (Ours)
Wikitext2 (ppl.) $\downarrow$ - $23.93$ $24.14$ - $9.44$ $10.10$ - $8.24$ $8.29$
HellaSwag (acc.) $\uparrow$ - $25.55\%$ $25.30\%$ - $82.10\%$ $81.01\%$ - $58.33\%$ 58.21%
GSM8k CoT (acc.) $\uparrow$ - - - - $72.84\%$ $70.58\%$ - $48.62\%$ $45.36\%$
@@
Table 1: Benchmarks: Perplexity & Accuracy for smolLM-1.7B-Instruct, phi3-mini-4k-Instruct and mistral-7B-Instruct on various language tasks
Property Models
smolLM-Instruct phi3-mini-4k-Instruct mistral-Instruct
Quantization Type 88 bits 88 bits 88 bits
Weight Compressibility SOTA Ours Improvement SOTA Ours Improvement SOTA Ours Improvement
Bits Saved $0.29$ $\mathbf{2.08}$ $\times 7.2$ $0.30$ $\mathbf{2.42}$ $\times 8.1$ $0.31$ $\mathbf{2.16}$ $\times 7.0$
Quantization Type 44 bits 44 bits 44 bits
Weight Compressibility SOTA Ours Improvement SOTA Ours Improvement SOTA Ours Improvement
Bits Saved $0.21$ $\mathbf{2.43}$ $\times 11.6$ $0.20$ $\mathbf{2.61}$ $\times 13.1$ $0.21$ $\mathbf{2.38}$ $\times 11.3$
Table 2: A comparison showing how our quantization scheme improves downstream entropy compressibility of weights versus SOTA quantization techniques
@@
Huffman weight encoding: Huffman encoding – which originates from entropy coding – is a popular data compression technique. The technique is based on the construction of a binary tree, where each symbol is assigned a score corresponding to its frequency of occurrence. This process involves recursively combining pairs of symbols into a tree structure. More frequent symbols are positioned closer to the root of the tree, resulting in shorter paths and consequently shorter binary codes for these symbols. The length of each symbol’s binary code is proportional to its distance from the root, thereby minimizing the total length of the encoded data. This ensures that no other prefix code can offer a more efficient representation of the data, making the Huffman encoding optimal for lossless compression. In this work, we employ Huffman encoding to compress the weights of quantized LLMs, enabling their deployment on edge devices. By applying Huffman encoding to the quantized weights of an LLM, particularly those with skewed value distributions, we achieve significant compression without sacrificing accuracy. As discussed in Section 2, a post-training quantized LLM can match the performance of a full-precision LLM for most language generation benchmarks. Huffman coding converts the model’s weight parameters into variable-length codes, assigning shorter codes to frequent weight values, which reduces both storage and bandwidth costs.
Segmentation for parallel Huffman decoding: To enable real-time inference with Huffman-encoded weights, we parallelize the inherently sequential decoding process. Standard Huffman decoding processes variable-length symbols one at a time, causing latency on edge devices. We address this by segmenting the encoded parameter space while preserving the original weight-tensor structure. Each segment has known boundaries and can be decoded independently, allowing parallel processing across threads without synchronization. To balance workloads, we shuffle multiple segments per thread, mitigating skewed symbol distributions and ensuring consistent decoding times. This approach removes the serial bottleneck, enabling fast, low-latency inference suitable for time-critical applications. Figure 3 illustrates the design.
4 Experiments